OpenStudy (anonymous):

how many different 6-digit numbers greater than 500 000 can be formed using 1 , 5 , 7 ,7 ,7 ,8

OpenStudy (anonymous):

u have the answer key..???

OpenStudy (anonymous):

100

OpenStudy (anonymous):

Yes...Got....it

OpenStudy (anonymous):

How many ways u can arrange 1 , 5 , 7 ,7 ,7 ,8

OpenStudy (anonymous):

6!/ 3!

OpenStudy (anonymous):

6!/3! = 6*5*4 = 120

OpenStudy (anonymous):

Nw .....it says that it should be greater than 500 000

OpenStudy (anonymous):

so....the numbers shuld not start with 1

OpenStudy (anonymous):

so...find the numebrs which start with 1

OpenStudy (anonymous):

5!/3! = 5*4 = 20

OpenStudy (anonymous):

so.......Total 120 - 20 = 100

OpenStudy (anonymous):

got it thanks.

OpenStudy (anonymous):

Welcome...